OpenStudy (anonymous):

Find the slope of the function at the given value: y=sin(2x) atx=−π

OpenStudy (ipwnbunnies):

Find the derivative of the given function. Plug in x = -π

OpenStudy (ipwnbunnies):

x = -pi .-.

OpenStudy (anonymous):

I did the problem, but I am not sure where I went wrong

OpenStudy (anonymous):

And we haven't gotten to derivatives yet

OpenStudy (anonymous):

We are on slopes of tangent and secant lines and stuff like that. I understand it is similar to derivatives, but we haven't actually gotten there yet. We need to use the average and instantaneous rates of change stuff

OpenStudy (ipwnbunnies):

This is asking for the slope at a specific point, which is instantaneous rate. Average rate would be the average rate of change between two points. I'm not sure how to calculate instantaneous rate at a point without the derivative, I think.

OpenStudy (anonymous):

Right, so thats the tangent line

OpenStudy (ipwnbunnies):

Ohh, yes, yes. The slope of the line tangent to the function at x = -pi

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Its a kutasoftware problem : http://cdn.kutasoftware.com/Worksheets/Calc/04%20-%20Slope%20at%20a%20Value.pdf

OpenStudy (ipwnbunnies):

Bleh. I wonder if there is a process I'm forgetting about.

OpenStudy (anonymous):

It involves

OpenStudy (anonymous):



OpenStudy (ipwnbunnies):

The slope of the tangent line is equivalent to the value of the derivative at the same point. Oh, ok, good. That's the limit definition of the derivative.

OpenStudy (anonymous):

Ok

OpenStudy (ipwnbunnies):

If you follow the definition of the derivative to your given function, applying the limit and such, you'll get the derivative. Then, you can plug in -pi for x.

OpenStudy (anonymous):

I did. Idk where I went wrong. I got 0 instead of 2. I was hoping someone could show me how to do it

OpenStudy (anonymous):

@girlover

OpenStudy (anonymous):

Ok, so my problem comes down to lim (h->0) sin(2x)/x

OpenStudy (anonymous):

How does that equal 2?

OpenStudy (anonymous):

Can someone show me how they solved it? I keep messing up

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Please, I have had this question up for like 2 hours

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (dangerousjesse):

I'm assuming you mean at x=\(-\pi\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I accidentally left out the space

OpenStudy (dangerousjesse):

No problem, let me work this out right quick.

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

Im going to go to bed soon, so please ASAP

OpenStudy (dangerousjesse):

\(\frac{ð (y = sin(2 x))}{ð x} ~where~ x = -\pi\) plug in \(\frac{ð (y = sin(2 \times-\pi))}{ð -\pi}\) \(y'(-\pi) = 2\)

OpenStudy (anonymous):

It is blurring out the symbols

OpenStudy (dangerousjesse):

Refresh the page

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i dont understand what you did. we have to use the instantaneous rate of change formula

OpenStudy (dangerousjesse):

That was the best I could find as an explanation, I know it's no help; it's better than nothing, I suppose.

OpenStudy (anonymous):

I appreciate it

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

lol 40 minutes until i can bump

OpenStudy (anonymous):

I have to go to bed, i will respond in the morning.